Loa is two years younger than three times Mary's age. If Mary's age in six years will be 3x-6, what is Loas's age now?
Answer by stanbon(48545)

Loa is two years younger than three times Mary's age. If Mary's age in six years will be 3x-6, what is Loas's age now?
---------------
Loa is two years younger than three times Mary's age. If Mary's age in six years will be 3x-6, what is Loas's age now?
Ages 6 years from now:
Mary: 3x-6
Loa: ?
----------
Ages now:
Mary:3x-12
Loa: ?-6
---------------
EQUATION:
?-6+2=3(3x-12)
?-4=9x-36
?=9x-32
-------------------
Loa's age now =?-6=9x-32-6=9x-38
Cheers,
Stan H.